INVESTIGATION OF THE B’ STATE OF CaF and CaCl

R. LAWRUSZCZUK, \underline{J. ROSTAS} and B. SOEP, Laboratoire de Photophysique Moléculaire, Université Paris Sud, 91405 Orsay Cédex, France.

The B’^{2}Δ states of CaF and CaCl have been excited directly, under specific experimental conditions, through the B’^{2}Δ-X^{2}Σ^{+} forbidden transition, and detected by collecting the A-X or B-X fluorescence. Rotational analysis of the B’-X(0-0) band provides accurate molecular parameters in good agreement with those obtained previously by Verges et al.^{a} in the case of CaF. For CaCl, the B’^{2}Δ state, observed for the first time lies at 19232 cm^{−1} above the ground state, in excellent agreement with its position (19319 cm^{−1}) predicted by the Ligand Field Theory. The spin-orbit coupling constant A^{SO} for CaF and CaCl are close, indicating that the molecular structure is mainly governed by the Ca^{+} metal ion and only slightly by the ligand. Several excitation and deexcitation mechanisms, which may contribute, will be discussed.

aVerges, Effantin, Bernard, Toupouzkhanian, Allouche, d’Incan and Barrow, J. Phys. B., At. Mol. Opt. Phys., 26, 279-284